A pharmacist receives a shipment of 17 bottles of a drug and has 4 of the bottles tested. If 6 of the 17 bottles are contaminated, what is the probability that less than 3 of the tested bottles are contaminated? Express your answer as a fraction or a decimal number rounded to four decimal places.
Answer #1

from hypergeometric distribution

probability that less than 3 of the tested bottles are contaminated

=P(X<=2)=P(X=0)+P(X=1)+P(X=2)=   =0.9013